The fraction which is not equal to 4/5 is (A) 40/50 (B) 12/15 (c) 16/20 (d) 9/15

Grade:12

1 Answers

Harshit Singh
3 years ago
Dear Student

All the options given in the question are further simplified as,
Option(A) 40/50 = 4/5
Option (B) 12/15
Divide both numerator and denominator by 3.
= 4/5
Option (C) 16/20
Divide both numerator and denominator by 4.
= 4/5
Option (D) 9/15
Divide both numerator and denominator by 3.
= 3/5
Therefore, 3/5≠4/5



Thanks